# Connectedness of the Gromov boundary of fine curve graphs

### v2

@@ Abstract
In this paper, we study the topological properties of the Gromov boundary of the fine curve graph of an orientable finite-type surface of genus at least 2. This graph consisting of topological curves has much richer dynamics than the classical curve graph. Using the techniques introduced by Wright [Wri23], we show that this boundary is (path) connected and that the spheres in non-separating fine curve graph are connected.
Keywords: fine curve graph, Gromov boundary, connectedness.
@@ 1 Introduction
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. In order to study the group of diffeomorphisms of surface $G$, Bowden, Hensel and Webb design for $S$ its fine curve graph $\mathcal{C}^{\dagger}(S)$ [BHW22], where the vertices are smooth curves instead of their isotopy classes, and the edges are defined by disjointness.
Indeed, the fine curve graph $\mathcal{C}^{\dagger}(S)$ shares a lot of similarity with the curve graph of $S$ and can be considered as an analogue of the curve graph for the mapping class group: the graph is Gromov hyperbolic [BHW22] and the group of homeomorphisms of surface $S$ is isomorphic to the automorphism group of the topological version of the fine curve graph [Lon+21].
But compared to the classical curve graph, the fine curve graph $\mathcal{C}^{\dagger}(S)$ is much more complicated and the dynamics of homeomorphism group on $\mathcal{C}^{\dagger}(S)$ is much richer than the mapping class group on the curve graph: $\mathcal{C}^{\dagger}(S)$ is a locally uncountable; there exist homeomorphisms on the surface $S$ that induce parabolic isometries on $\mathcal{C}^{\dagger}(S)$ and by consequences there are points on the Gromov boundary that are not represented by ending laminations [BHW22, Bow+22], whereas it is the case for curve graphs [Kla22].
The study of geometric or topological interpretation of points at infinity in the fine curve graph is still an ongoing project. Efforts are made in understanding the stabilisers of these points in terms of (homological) rotation sets [Bow+22, GM23a, GM23].
In this paper, we are more interested in the topological properties of the Gromov boundary of fine curve graphs. This paper is written as one of the initiatives in this direction.
For the curve graph, Peter Storm asked an interesting question about the (path)-connectedness of its the Gromov boundary, which is recorded in [KL08, Question 10]. This question is answered positively by [Gab09, LS09, LMS11]. Recently, [Wri23] provides a new proof and proves that the boundary of curve graph is in addition linearly connected (see Definition 2.10).
Let $\operatorname{\mathcal{NC}}^{\dagger}(S)$ be the subgraph of $\mathcal{C}^{\dagger}(S)$ consisting of non-separating curves.
In this paper, we conclude the following results for the topology of Gromov boundary of the fine curve graph:
@@
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. Then the Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ of the fine curve graph is path connected.
Moreover, for any $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ and any small enough $\varepsilon>0$, if $b\in(1,1+\varepsilon)$ and $\rho_{o,b}$ is the visual metric defined on the Gromov boundary, then it is linearly connected.
We remark that Theorem 1.1 cannot be deduced from the path-connectedness of the Gromov boundary of curve graphs, since there are boundary points in the fine curve graph that are cannot be represented by ending laminations and that path-connectedness is usually not preserved under taking closures.
We say that a subgraph $G$ of a graph $\Gamma$ is connected if for any two points $a,b\in G$, there exists a finite sequence $a=v_{0},\ v_{1},\ \dots,\ v_{n}=b$ such that $v_{i}$ is connected to $v_{i+1}$ by an edge for all $10\leq i\leq n-$ 1 and $v_{i}\in G$ for all $0\leq i\leq n$.
@@
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. For any integer $r>0$ and any $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$, the $r$-sphere $\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,o)=r\}$ is connected.
The main difficulties of dealing with the fine curve graph is the non-transversality between topological curves and the lack of analogue in the setting of fine curve graphs of tight geodesics as in the curve graphs. In order to deduce the results, we first generalise the criterion introduced by Wright in [Wri23] to any hyperbolic graphs with no “dead ends” (see Proposition 2.11). In order to apply Proposition 2.11, it suffices to check that the assumptions are satisfied by a subgraph quasi-isometric to $\mathcal{C}^{\dagger}(S)$. In principle, under the non-separating and transverse settings, as long as there are only finitely many topological curves involved, the fine curve graph looks like the curve graph of a punctured subsurface (see Proposition 3.5), which allows us to make use of [Wri23, Proposition 5.4]. The technical part of the proof lies in passing non-transverse pairs of curves into transverse ones. At the end of this paper, we also gave several interesting observations on the group action of $\operatorname{Homeo}(S)$ on the Gromov bordification of $\mathcal{C}^{\dagger}(S)$.
@@ Acknowledgement
The first author wishes to thank the hospitality of Guangxi Center for Mathematical Research for a visit, during which this work is partially carried out. The second author also thanks the hospitality of Tsinghua University and Sun Yat-sen University. The first author is supported by CSC-202108170018. The second author is supported by the National Natural Science Foundation of China (No. 12001122, 12271533, 12361014) and the special grants for Guangxi Ba Gui Scholars. The authors are grateful to Yongsheng Jia for informing discussion and to Alex Wright for explaining to us some detail in his work. The authors also thank Bruno Duchesne and Pierre-Antoine Guihéneuf for reviewing the early drafts of this paper.
@@ 2 Gromov hyperbolic space and bordification
The notion of $\delta$-hyperbolic space is originally introduced in [Gro87]. Classical references are [BH13] and [DSU17], while the latter treats more general cases. Here we cover a very limited part of this topic to make clarified some folklore results and introduce the tools that we will later need. For more details, we refer the reader, if interested, to the classical references above.
@@
$\langle x,y\rangle_{o}\coloneqq\frac{1}{2}\big{(}d_{X}(x,o)+d_{X}(y,o)-d_{X}(x,y)\big{)}$ (1)
for every $x,y\in X$. The quantity $\langle x,y\rangle_{o}$ in (1) is called the Gromov product of $x$ and $y$ based at $o$, or simply the Gromov product of $x$ and $y$.
@@
$\langle x,y\rangle_{o}\geq\min\left\{\langle x,z\rangle_{o},\langle z,y\rangle_{o}\right\}-\delta$ (2)
@@
When the metric space is geodesic, i.e. a space where every pair of points can be connected by a path isometric to a closed real interval, $\delta$-hyperbolicity is equivalent to the property that every geodesic triangle is $\delta^{\prime}$-slim, for some uniform $\delta^{\prime}\geq 0$. This condition is attributed to Rips by Gromov in [Gro87] and is thus also often called Rips condition, see [DSU17, §4.3] for the exact definition.
@@
Let $(X,d_{X})$ be a Gromov hyperbolic space. Then we define its Gromov boundary by the equivalent classes of Cauchy-Gromov sequences in $X$ and we denote it by $\partial X$. We say that a sequence $(x_{n})_{n\geq 0}$ converges to a boundary point $\xi\in\partial X$ if it is a representative of $\xi$. The disjoint union $\mathrm{bord}X\coloneqq X\cup\partial X$ is called the Gromov bordification of $X$.
@@
$\left|\lim_{n,m\to\infty}\langle x_{n},y_{m}\rangle_{o}-\langle\xi,\eta\rangle_{o}\right|=O(\delta)$ (3)
for any Cauchy-Gromov sequences $(x_{n})\in\xi$ and $(y_{n})\in\eta$ and any $o\in X$ [DSU17, Lemma 3.4.7].
With Gromov product, the Gromov boundary can be endowed with the uniform structure generated by the entourages in the form of
@@
$\left\{\xi\in\partial X\colon\langle x,\xi\rangle_{o}>R>0\right\}.$ (4)
This fact can be easily verified by (2) and (3) under the settings of Gromov hyperbolicity.
Recall that a map $f\colon(X,d_{X})\to(Y,d_{Y})$ between two metric spaces is a $(\lambda,k)$-quasi-isometric embedding if there exists $\lambda\geq 1$ and $k>0$ such that
@@
for every $x,y\in X$. If in addition, the map $f$ is essentially surjective, i.e. there exists $C>0$ such that $d_{Y}\big{(}y,f(X)\big{)}<C<\infty$ for all $y\in Y$, then $f$ is a quasi-isometry and we say that $X$ and $Y$ are quasi-isometric. We remark that being quasi-isometric is an equivalent relation. A quasi-geodesic in a metric space $X$ is a quasi-isometric embedding of a real interval into $X$.
The following result is pivotal in large scale geometry à la Gromov and is well-known in proper cases. But we remark that it actually holds for non-proper spaces [Has22, Väi05]. For convenience reason, we state it here below. There are several different proofs from different approaches and the one we present here might have already been known to [Cap+15]:
@@
Let $(X,d_{X})$ be a $\delta$-hyperbolic geodesic space and $o\in X$ be an arbitrary base point. Then $\partial X$ is in bijection with the equivalent classes of quasi-geodesic rays issued from $o$ and two quasi-geodesic rays are equivalent if they have bounded Hausdorff distance. Moreover, a quasi-isometric embedding $f\colon X\to Y$ induces a topological embedding $f_{\partial}\colon\partial X\to\partial Y$ and $f_{\partial}$ is a homeomorphism if $f$ is essentially surjective.
@@
In [Gro87, §7], a Gromov hyperbolic space $X$ is ultracomplete if for any two points $x,y\in\mathrm{bord}X$, there exists a geodesic connecting $x$ to $y$. The proof above actually shows that $u(X)$ is ultracomplete. If a Gromov hyperbolic geodesic space is proper, then it is ultracomplete; but the converse is not true: the curve graph of a compact hyperbolic surface is ultracomplete but not proper [Min10, Lemma 5.14].
Another powerful tool to study the Gromov boundary is the following notion of visual metric that is carefully treated in [DSU17, §3.6] and [Väi05] under general settings:
@@
Let $(X,d_{X})$ be a Gromov hyperbolic space. Then $\partial X$ is completely metrizable. For any based point $o\in X$, the topology is compatible to the visual metric $\rho_{o,b}$ from $o\in X$ with parameter $b\in(1,1+\varepsilon)$ for some small enough $\varepsilon>0$. Moreover, we have the estimation
$b^{-\langle\xi,\eta\rangle_{o}}/4\leq\rho_{o,b}(\xi,\eta)\leq b^{-\langle\xi,\eta\rangle_{o}}$ (5)
@@
The first result about this notion is referred as the reverse triangle inequality:
Proposition 2.6 ([MT18], Proposition 2.2).
@@
Clearly from (6), the nearest point projection is coarsely well-defined: the diameter of $\pi_{\gamma}(X)$ is bounded uniformly by a multiple of $\delta$.
Using this inequality, we can deduce the following estimation, which is essentially a translation into nearest points projection language of the thinness of geodesic triangles in Gromov hyperbolic spaces:
@@
$\langle x,y\rangle_{o}=d_{X}(o,p)+O(\delta).$ (7)
@@
Using (6) for $d_{X}(x,o)$ and $d_{X}(x,y)$ appeared in (1) will yield the result. ∎
@@
By continuity of $\ell_{1}$ and $\ell_{2}$, if $x\coloneqq\ell_{1}(r)$ and $y\coloneqq\ell_{2}(r)$, then $d_{X}(x,y)=E$. Let $\gamma\coloneqq[o,y]$ be a geodesic segment and $p\in\pi_{\gamma}(x)$. Applying the reverse triangle inequality (6) twice, we have
@@
$\frac{r}{\lambda}-k-E/2+O(\delta)\leq d_{X}(o,p)\leq\lambda r+k-E/2+O(\delta).$ (9)
But by (3), when $E$ is sufficiently large, we have $\langle x,y\rangle_{o}=\langle\xi,\eta\rangle_{o}+O(\delta)$. Combining this fact with (7) and (9), we can deduce that
@@
By setting $C\coloneqq 4b^{k+E/2+O(\delta)}$, we can deduce the desired inequality via applying the estimation (5). ∎
@@
The application of Lemma 2.8 is the following proposition inspired by [MS20, Proof of Proposition 5.2] and is an improvement of [Wri23, Proposition 2.2].
@@
The proof is a straightforward adaption of the proof of [Wri23, Proposition 2.2]. Let $\xi_{0},\xi_{1}\in\partial X$ be any two points on the Gromov boundary. Now connect $\xi_{0}$ and $\xi_{1}$ to $o$ respectively by continuous $(1,k)$-quasi-geodesic rays $\gamma_{0}$ and $\gamma_{1}$. By the same arguments in [Wri23, Proof of Proposition 2.2], (G1) assures the existence of a dense subset $I\subset[0,1]$ such that for all $t\in I$, there is a geodesic ray $\gamma_{t}\subset\Gamma$ connecting $o$ to a point $c(t)\in\partial\Gamma$. Since the family $\left(\gamma_{t}\right)_{t\in I\cup\{0,1\}}$ consists of continuous $(1,k)$-quasi-geodesic rays, thus by Lemma 2.8, we can deduce from (G2) and (G3) the same estimation as in [Wri23, Proof of Proposition 2.2], i.e. there exists $L>0$ independent of $\xi_{0}$ and $\xi_{1}$ such that the diameter $\mathrm{diam}\big{(}c(I)\big{)}<L\rho_{o,b}(\xi_{0},\xi_{1})$ and that $c\colon I\to\partial\Gamma$ is uniformly continuous. The desired path is then the closure of $c(I)$. ∎
@@
Proposition 2.11 cannot be obtained by applying [Wri23, Proposition 2.2].
3 Fine curve graph
Let $S$ be an orientable finite-type surface of genus $g\geq 2$. It can carry a hyperbolic metric. In a recent paper, Bowden, Hensel and Webb design for $S$ its fine curve graph [BHW22] as a combinatorial tool for studying $\operatorname{Homeo}(S)$, which is an analogue to the renown curve graph.
An essential simple closed smooth curve $\gamma$ is a proper $\mathcal{C}^{\infty}$-embedding of the circle $S^{1}\hookrightarrow S$ such that $\gamma$ is not homotopic to a singleton nor bounds a once-punctured disc or an annulus. If in addition $S-\gamma$ is connected, then the curve is non-separating.

### v4

# biblio.bib

@@ Abstract
The fine curve graph was introduced to study homeomorphism group of surfaces. In this paper we study the topology of the Gromov boundary of this graph for closed surfaces with higher genus. We first prove a bounded geodesic image theorem for the fine curve graph, a consequence of which is the non-compactness of the Gromov boundary. Using this theorem, we are able to show that the Gromov boundary is linearly connected with respect to some visual metric.
Keywords: fine curve graph, bounded geodesic image theorem, Gromov boundary, connectedness.
@@ 1 Introduction
Let $S$ be an orientable connected closed surface. In their work on the diffeomorphism groups of surfaces, Bowden, Hensel, and Webb introduced the concept of the fine curve graph, denoted by $\mathcal{C}^{\dagger}(S)$ [bowden2022quasi]. This graph differs from the classical curve graph in that its vertices correspond directly to simple closed essential curves on $S$, rather than the isotopy classes of these curves. An edge in $\mathcal{C}^{\dagger}(S)$ connects two vertices if the corresponding curves are disjoint. This construction enables a finer analysis of the geometry and dynamics associated with diffeomorphisms on $S$.
Indeed, the fine curve graph $\mathcal{C}^{\dagger}(S)$ shares many similarities with the classical curve graph of $S$ and serves as an analogue of the curve graph for the mapping class group: the graph $\mathcal{C}^{\dagger}(S)$ is also Gromov hyperbolic [bowden2022quasi], and the group of homeomorphisms of the surface $S$ is isomorphic to the automorphism group of the fine curve graph [long2021automorphisms].
Yet compared to the classical curve graph, the fine curve graph $\mathcal{C}^{\dagger}(S)$ is significantly more complex, and the dynamics of the homeomorphism group acting on $\mathcal{C}^{\dagger}(S)$ are much richer than those of the mapping class group on the curve graph. Notably, $\mathcal{C}^{\dagger}(S)$ is locally uncountable. Although there are points on the Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ that can be represented by certain foliations on the surface [bowden2024boundaryfinecurvegraph], there also exist homeomorphisms of the surface $S$ that induce parabolic isometries on $\mathcal{C}^{\dagger}(S)$, which in turn leads to boundary points not representable by foliations [bowden2022quasi, bowden2022rotation], in contrast to the boundary of the curve graph, where all points are represented by minimal measurable foliations [klarreich2022boundary].
The geometric and topological interpretation of points at infinity in the fine curve graph remains an active area of research. Current efforts aim to understand the stabilisers of these points through (homological) rotation sets [bowden2022rotation, guiheneuf2023parabolic, guiheneuf2023hyperbolic] and foliations [bowden2024boundaryfinecurvegraph].
In this paper, our focus is on exploring the topological properties of the Gromov boundary of the fine curve graph. This work represents one of the initial steps in this direction.
For the curve graph, Peter Storm posed an interesting question about the (path)-connectedness of its Gromov boundary, as recorded in [kent2008shadows, Question 10]. This question was answered affirmatively in [gabai2009almost, leininger2009connectivity, leininger2011universal]. Recently, [wright2023spheres] provided a new proof, showing additionally that the Gromov boundary of the curve graph is linearly connected (see Definition 2.10).
For any $o\in\mathcal{C}^{\dagger}(S)$ and sufficiently small $\varepsilon>0$, a visual metric $\rho_{o,b}$ can be defined on the Gromov boundary, where $b\in(1,1+\varepsilon)$, and this metric is compatible with the topology on the Gromov boundary.
The main goal of this paper is to show the following result of linear connectedness with respect to some visual metric:
@@
Let $S$ be an orientable connected closed surface with genus $g\geq 3$, then the Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ of the fine curve graph is path connected.
Moreover, if $o$ is a non-separating curve on $S$, then equipped with the visual metric $\rho_{o,b}$, the Gromov boundary is linearly connected.
Following the approach in [wright2023spheres], this result is obtained by analysing a subgraph that is quasi-isometric to the original Gromov hyperbolic graph. This technique originates in [Bestvina1991TheBO].
In Section 2, we first generalise the criterion introduced by Wright in [wright2023spheres] to apply to any hyperbolic graph without “dead ends” (see Proposition 2.11). We then apply Proposition 2.11 to a carefully selected subgraph.
In our setting, this subgraph is $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, the subgraph of $\mathcal{C}^{\dagger}(S)$ consisting of non-separating curves that are transverse to $o$. This subgraph is quasi-isometric to the subgraph $\mathcal{NC}^{\dagger}(S)$ of non-separating curves, which in turn is quasi-isometric to the fine curve graph $\mathcal{C}^{\dagger}(S)$. In addition, the spheres in the subgraph $\mathcal{NC}^{\dagger}(S)$ are connected.
@@
Let $S$ be an orientable connected closed surface with genus $g\geq 3$. For any integer $r>0$ and any $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$, the $r$-sphere $\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,o)=r\}$ is connected.
However, the connectedness of spheres in the curve graph and in its Gromov boundary does not directly help in proving Theorem 1.1. Since the fine curve graph is approximated by the surviving curve graph on punctured subsurfaces $S-P$ (see Proposition 3.7), the paths connecting two points on the sphere obtained by the method in [wright2023spheres] may involve pant curves on the punctured subsurfaces, which are not essential curves on $S$.
To address the complications introduced by pant curves, we establish a bounded geodesic image theorem for the fine curve graph in Section 4, following the approach of [webb2013shortproofboundedgeodesic].
Theorem 1.3 (Bounded image theorem).
Given a surface $S$ with $g(S)\geq 2$, there exists a constant $M$ such that whenever $Y$ is a subsurface and $g=\left(\gamma_{i}\right)$ is a geodesic in $\mathcal{C}^{\dagger}(S)$ such that $\gamma_{i}$ intersects $Y$ essentially for all $i$, then $\operatorname{diam}_{d_{Y}^{\dagger}}(\pi_{Y}(g))\leq M$.
Similarly, we also have a version for the non-surviving curve graph (see Theorem 4.12), which is the theorem used to prove the connectedness properties. The primary challenges in the proof arise from the non-transversality between topological curves and the surgeries on the surviving curve graph, both of which are carefully treated in Sections 3 and 4.
With the bounded geodesic image theorem, we can also conclude that the Gromov boundary of the fine curve graph is not compact, similar to the case of the curve graph.
Theorem 1.4.
Let $S$ be an orientable connected closed surface with genus $g(S)\geq 2$. Then the Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ is not compact.
To verify the conditions in Proposition 2.11, the key step is to use the bounded geodesic image theorem to construct paths connecting points on thickened spheres in the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, as desired. These constructions are addressed in Section 5.
At the end of this paper, we also present several interesting observations concerning the group action of $\operatorname{Homeo}(S)$ on the Gromov bordification of $\mathcal{C}^{\dagger}(S)$.
@@ Acknowledgement
The first author wishes to thank the hospitality of Guangxi Center for Mathematical Research. The first author is supported by CSC-202108170018 and Tremplin–ERC Starting Grant MAGIC (reference ANR-23-TERC-0007). The second author is partially supported by National Natural Science Foundation of China grant No.12001122, 12271533, 12361014. The authors are grateful to Mladen Bestvina, Bruno Duchesne, Federica Fanoni, Pierre-Antoine Guihéneuf, Sebastian Hensel, Yongsheng Jia, Roberta Shapiro and Alex Wright for their valuable comments.
@@ 2 Gromov hyperbolic space and bordification
2.A. Definition.
The notion of $\delta$-hyperbolic space was originally introduced in [gromov1987hyperbolic]. Classical references are [bridson2013metric] and [das2017geometry], while the latter treats more general cases. Here we cover a very limited part of this topic to clarify some folklore results and introduce the tools that we will later need. For more details, we refer the reader, if interested, to the classical references above.
@@
$\langle x,y\rangle_{o}\coloneqq\frac{1}{2}\big{(}d_{X}(x,o)+d_{X}(y,o)-d_{X}(x,y)\big{)}$ (2.1)
for every $x,y\in X$. The quantity $\langle x,y\rangle_{o}$ in (2.1) is called the Gromov product of $x$ and $y$ based at $o$, or simply the Gromov product of $x$ and $y$.
@@
$\langle x,y\rangle_{o}\geq\min\left\{\langle x,z\rangle_{o},\langle z,y\rangle_{o}\right\}-\delta$ (2.2)
@@
When the metric space is geodesic, i.e. a space where every pair of points can be connected by a path isometric to a closed real interval, $\delta$-hyperbolicity is equivalent to the property that every geodesic triangle is $\delta^{\prime}$-slim, for some uniform $\delta^{\prime}\geq 0$. This condition is attributed to Rips by Gromov in [gromov1987hyperbolic] and is thus also often called Rips condition, see [das2017geometry, §4.3] for the exact definition.
@@
Let $(X,d_{X})$ be a Gromov hyperbolic space. Then we define its Gromov boundary by the equivalence classes of Cauchy-Gromov sequences in $X$ and we denote it by $\partial X$. We say that a sequence $(x_{n})_{n\geq 0}$ converges to a boundary point $\xi\in\partial X$ if it is a representative of $\xi$. The disjoint union
$\mathrm{bord}X\coloneqq X\cup\partial X$
is called the Gromov bordification of $X$.
@@
$\left|\lim_{n,m\to\infty}\langle x_{n},y_{m}\rangle_{o}-\langle\xi,\eta\rangle_{o}\right|=O(\delta)$ (2.3)
for any Cauchy-Gromov sequences $(x_{n})\in\xi$ and $(y_{n})\in\eta$ and any $o\in X$ [das2017geometry, Lemma 3.4.7].
With the Gromov product, the Gromov boundary can be endowed with the uniform structure generated by the entourages in the form of
@@
$\left\{\xi\in\partial X\colon\langle x,\xi\rangle_{o}>R>0\right\}.$ (2.4)
This fact can be easily verified by (2.2) and (2.3) under the setting of Gromov hyperbolicity.
Recall that a map $f\colon(X,d_{X})\to(Y,d_{Y})$ between two metric spaces is a $(\lambda,k)$-quasi-isometric embedding if there exist $\lambda\geq 1$ and $k>0$ such that
@@
for every $x,y\in X$. If in addition, the map $f$ is essentially surjective, i.e. there exists $C>0$ such that $d_{Y}\big{(}y,f(X)\big{)}<C<\infty$ for all $y\in Y$, then $f$ is a quasi-isometry and $X$ and $Y$ are said to be quasi-isometric. We remark that being quasi-isometric is an equivalence relation. A quasi-geodesic in a metric space $X$ is a quasi-isometric embedding of a real interval into $X$.
The following result is pivotal in large scale geometry à la Gromov and is well-known in proper cases. But we remark that it actually holds for non-proper spaces [hasegawa2022gromov, vasala2005gromov]. For convenience, we state it here below. There are several different proofs from different approaches and the one we present here might already be known to [caprace2015amenable]:
@@
Let $(X,d_{X})$ be a $\delta$-hyperbolic geodesic space and $o\in X$ be an arbitrary base point. Then $\partial X$ is in bijection with the equivalence classes of quasi-geodesic rays issued from $o$ and two quasi-geodesic rays are equivalent if they have bounded Hausdorff distance. Moreover, a quasi-isometric embedding $f\colon X\to Y$ induces a topological embedding $f_{\partial}\colon\partial X\to\partial Y$ and $f_{\partial}$ is a homeomorphism if $f$ is essentially surjective.
@@
In [gromov1987hyperbolic, §7], a Gromov hyperbolic space $X$ is ultracomplete if for any two points $x,y\in\mathrm{bord}X$, there exists a geodesic connecting $x$ to $y$. The proof above actually shows that $u(X)$ is ultracomplete. If a Gromov hyperbolic geodesic space is proper, then it is ultracomplete. But the converse is not true: the curve graph of a compact hyperbolic surface is ultracomplete but not proper [minsky2010classification, Lemma 5.14].
2.B. Visual metric and boundary connectedness.
Another powerful tool to study the Gromov boundary is the following notion of visual metric that is carefully treated in [das2017geometry, §3.6] and [vasala2005gromov] under general settings:
@@
Let $(X,d_{X})$ be a Gromov hyperbolic space. Then $\partial X$ is completely metrizable. For any based point $o\in X$, the topology is compatible to the visual metric $\rho_{o,b}$ from $o\in X$ with parameter $b\in(1,1+\varepsilon)$ for some small enough $\varepsilon>0$. Moreover, we have the following estimate
$b^{-\langle\xi,\eta\rangle_{o}}/4\leq\rho_{o,b}(\xi,\eta)\leq b^{-\langle\xi,\eta\rangle_{o}}$ (2.5)
@@
The first result about this notion is referred to as the reverse triangle inequality:
Proposition 2.6 ([maher2018random], Proposition 2.2).
@@
Clearly from (2.6), the nearest point projection is coarsely well-defined: the diameter of $\pi_{\gamma}(X)$ is bounded uniformly by a multiple of $\delta$.
Using this inequality, we can deduce the following estimate, which essentially translates the thinness of geodesic triangles in Gromov hyperbolic spaces into the language of nearest point projection:
@@
$\langle x,y\rangle_{o}=d_{X}(o,p)+O(\delta).$ (2.7)
@@
Applying (2.6) for $d_{X}(x,o)$ and $d_{X}(x,y)$ appeared in (2.1) yields the result. ∎
@@
By continuity of $\ell_{1}$ and $\ell_{2}$, if $x\coloneqq\ell_{1}(r)$ and $y\coloneqq\ell_{2}(r)$, then $d_{X}(x,y)=E$. Let $\gamma\coloneqq[o,y]$ be a geodesic segment and $p\in\pi_{\gamma}(x)$. Applying the reverse triangle inequality (2.6) twice, we have
@@
$\frac{r}{\lambda}-k-E/2+O(\delta)\leq d_{X}(o,p)\leq\lambda r+k-E/2+O(\delta).$ (2.9)
But by (2.3), when $E$ is sufficiently large, we have $\langle x,y\rangle_{o}=\langle\xi,\eta\rangle_{o}+O(\delta)$. Combining this fact with (2.7) and (2.9), we can deduce that
@@
By setting $C\coloneqq 4b^{k+E/2+O(\delta)}$, we can deduce the desired inequality via applying the estimate (2.5). ∎
@@
The application of Lemma 2.8 is the following proposition inspired by [mackey2020quasi, Proof of Proposition 5.2] and is an improvement of [wright2023spheres, Proposition 2.2].
@@
The proof is a straightforward adaptation of the proof of [wright2023spheres, Proposition 2.2]. Let $\xi_{0},\xi_{1}\in\partial X$ be any two points on the Gromov boundary. Now connect $\xi_{0}$ and $\xi_{1}$ to $o$ respectively by continuous $(1,k)$-quasi-geodesic rays $\gamma_{0}$ and $\gamma_{1}$. By the same arguments in [wright2023spheres, Proof of Proposition 2.2], (G1) ensures the existence of a dense subset $I\subset[0,1]$ such that for all $t\in I$, there is a geodesic ray $\gamma_{t}\subset\Gamma$ connecting $o$ to a point $c(t)\in\partial\Gamma$. Since the family $\left(\gamma_{t}\right)_{t\in I\cup\{0,1\}}$ consists of continuous $(1,k)$-quasi-geodesic rays, by Lemma 2.8, we can deduce from (G2) and (G3) the same estimation as in [wright2023spheres, Proof of Proposition 2.2], i.e. there exists $L>0$ independent of $\xi_{0}$ and $\xi_{1}$ such that the diameter $\mathrm{diam}\big{(}c(I)\big{)}<L\rho_{o,b}(\xi_{0},\xi_{1})$ and that $c\colon I\to\partial\Gamma$ is uniformly continuous. The desired path is then the closure of $c(I)$. ∎
@@
Proposition 2.11 cannot be obtained by applying [wright2023spheres, Proposition 2.2].
3 Fine curve graph and its subgraphs
3.A. Fine curve graph.
Let $S$ be an orientable connected closed surface of finite type with genus $g\geq 2$. Such a surface admits a complete hyperbolic metric. In a recent paper, Bowden, Hensel, and Webb introduced the fine curve graph of $S$ [bowden2022quasi], a combinatorial tool designed to study the homeomorphism group $\operatorname{Homeo}(S)$ and the diffeomorphism group $\mathrm{Diff}(S)$ of $S$. This construction serves as an analogue to the well-known curve graph.
An essential simple closed smooth curve $\gamma$ on $S$ is a smooth embedding of the circle, $\gamma:S^{1}\hookrightarrow S$, with the property that $\gamma$ does not bound a disc. Unless stated otherwise, all curves considered on $S$ will be assumed to be essential, simple, closed, and smooth.
The fine curve graph $\mathcal{C}^{\dagger}(S)$ is a graph whose vertices correspond to essential simple closed smooth curves on $S$. Two vertices are connected by an edge if the corresponding curves are disjoint. Equipped with the simplicial metric, the fine curve graph is Gromov hyperbolic [bowden2022quasi].